I need help with this word problem. A herd of an endangered antelope species was relocated to a wildlife refuge. The exponential model for the antelope population after t years is given by the equation P=3000/(1+9e^-0.18310t)
where P represents the population after t years. Find the population after 6 years.

Answer by tutorcecilia(2144)

P=3000/(1+9e^-0.18310t)
P(6)=3000/(1+9e^-0.18310(6)) [plug-in t=6 and solve]
P(6) = 749.99 antelope after 6 years.